# TLOB: A Novel Transformer Model with Dual Attention for Stock Price Trend Prediction with Limit Order Book Data

### v2

@@ 5.1. MLPLOB
(8) $\mathbf{U}_{i,*}\;=\;\sigma\Bigl(\text{LayerNorm}\bigl(\sigma(\mathbf{X}_{i,*}\,\mathbf{W}_{1})\,\mathbf{W}_{2}\;+\;\mathbf{X}_{i,*}\bigr)\Bigr)\quad\text{for }i=1,\dots,T,$
where $\sigma$ is the GeLU activation function (Hendrycks and Gimpel, 2016), and LayerNorm denotes layer normalization.
@@
(9) $\mathbf{Z}_{*,j}\;=\;\sigma\Bigl(\text{LayerNorm}\bigl(\sigma(\mathbf{U}_{*,j}\,\mathbf{W}_{3})\,\mathbf{W}_{4}\;+\;\mathbf{U}_{*,j}\bigr)\Bigr)\quad\text{for }j=1,\dots,N.$
@@
Final Prediction. After several blocks of feature and temporal mixing, MLPLOB performs dimensionality reduction to collapse all features into a single vector, which then passes through several fully connected layers that gradually diminish the vector dimension and a final standard classification head. The network outputs the directional trend (up, down, or stable) for the final time step. Our primary objective in devising MLPLOB is to show that a carefully structured MLP-based model can match or exceed more complex architectures in the SPTP task. The same method is also applied to TLOB.
5.2. TLOB
The Transformer architecture (Vaswani et al., 2017) has led to major breakthroughs in deep learning, notably in natural language processing (Brown, 2020; Khan et al., 2022) and time-series modeling (Wen et al., 2022). A key advantage is the ability to capture long-range dependencies without suffering as much from vanishing gradients or forgetting, and performance typically scales favorably with increased data (Kaplan et al., 2020). Because massive volumes of financial data (in the terabyte range) are available, Transformers are well-positioned for LOB modeling.
@@ 5.2. TLOB
Temporal vs. Feature Attention. While standard Transformers (Vaswani et al., 2017) process tokens along a single dimension, LOB data naturally requires both temporal and spatial dependencies to be learned (Sirignano, 2019; Sirignano and Cont, 2021). For instance, time-step $t$ can reveal how deeper or shallower levels relate to one another, as well as how trends evolve over past snapshots. Hence, dual-attention explicitly addresses these two axes of variation. To investigate the importance of each type of attention layers we performed an ablation study (Section 7.5).
Bilinear Normalization Layer. To address non-stationarity and magnitude disparity (prices and sizes) in financial time series, we employ a Bilinear Normalization layer (Tran et al., 2021) as the initial layer. Unlike conventional $z$-score normalization, which can fail under distribution shifts at inference time, bilinear normalization adapts to batch-specific statistics, maintaining robust performance even when market conditions change. The same layer is also used in MLPLOB.
Positional Encoding. Because self-attention is permutation-invariant, we incorporate sinusoidal positional embeddings (Vaswani et al., 2017) to preserve the chronological structure within each LOB window. This embedding ensures that TLOB respects the temporal ordering of snapshots, which is crucial for modeling price evolution.
@@
6. Experiments
We conduct a comprehensive evaluation of MLPLOB and TLOB model training and testing on both the Benchmark FI-2010 dataset and the TSLA-INTC dataset, composed of Tesla and Intel. TLOB and MLPLOB surpass SoTA performances on every dataset and every horizon. TLOB performs the best on larger horizons, while MLPLOB performs the best on the shorter ones. Our experiments extend beyond merely demonstrating the state-of-the-art performance of TLOB, aiming to address several critical research questions: (1) Are stock prices harder to forecast than in the past? (2) What if we choose $\theta$ equal to the average spread? (3) Are temporal and spatial attention necessary? Through these investigations, we seek not only to validate our models’ predictive capabilities but also to contribute to the broader understanding of deep learning applications in financial forecasting.
Table 1. Intel and Tesla main features for January 2015. Average liquidity is computed as the average quantity available in the first 10 LOB levels.
@@ 6. Experiments
TSLA $0.42\pm 2.84$ $23,927,602\pm 4,554,884$ $0.16$ $3,320$
INTC $0.44\pm 1.66$ $304,325,400\pm 69,340,430$ $0.01$ $124,960$
6.1. TSLA-INTC Dataset
In the majority of state-of-the-art (SoTA) research within the domain of Deep Learning applied to LOB data, researchers typically employ one, two, or three stocks (Shi and Cartlidge, 2023; Coletta et al., 2021; Nagy et al., 2023; Li et al., 2020; Shi and Cartlidge, 2022; Hultin et al., 2023), predominantly from the technology sector. Adhering to this established practice, we construct a LOB dataset comprising two NASDAQ-listed stocks namely, Tesla and Intel – spanning the period from January 2nd to January 30th, 2015. We posit that stylized facts and market microstructure characteristics exhibit independence from individual stock behaviors (as demonstrated in (Bouchaud et al., 2009, 2002; Cont et al., 2014; Gould et al., 2013) [^3] These seminal works in finance elucidate the universal statistical properties of LOBs, transcending specific stocks and markets.), thereby rendering specific stock attributes non-critical to the analysis. The dataset encompasses 20 order book files for each stock, corresponding to each trading day, resulting in a total of approximately 24 million samples. Each order book sample is represented as a tuple $\big(P^{ask}(t),V^{ask}(t),P^{bid}(t),V^{bid}(t)\big)$, where $P^{ask}(t)$ and $P^{bid}(t)\in\mathbb{R}^{L}$ denote the prices at levels $1$ through $L$, and $V^{ask}(t)$ and $V^{bid}(t)\in\mathbb{R}^{L}$ represent the corresponding volumes. The dataset is partitioned such that the initial 17 days are allocated for training, the 18th day for validation, and the final two days for testing. In Table 1 we present the main features of Tesla and Intel for January 2015. In Figure 3 we show the mid-price traces. As shown both the features and the mid-price traces are very different, making the evaluation more general. Unfortunately, we cannot make the dataset public for copyright reasons.
Sampling. Limit Order Book data, especially for liquid stocks, is massive, every day, hundreds of thousands of orders are placed for each stock. Furthermore, financial data are known to have a low signal-to-noise ratio (Nagel, 2021). Accordingly, it is unnecessary to consider every LOB update, so defining a valid sampling technique is essential. While time-based and event-based sampling methods  [^4] FI-2010 uses this type of sampling, a LOB snapshot is taken every 10 events. are used, they fail to capture the varying impact of transactions. In fact, single transactions can have very different impacts on the market. Volume-based sampling offers a solution by sampling the LOB after a predetermined volume of shares has been traded, thus reflecting the magnitude of market activity. Therefore, we adopted a sampling strategy based on trading volume, where snapshots of the Limit Order Book (LOB) are taken every 500 stocks traded. This method achieves a compromise between maintaining adequate temporal consistency within windows and ensuring significant variation between samples.
Figure 3. Tesla and Intel mid-price traces for January 2015.
6.2. Benchmark dataset FI-2010
@@
Our model will be evaluated against SoTA models utilizing the FI-2010 benchmark dataset (Ntakaris et al., [n. d.]). The FI-2010 dataset (Ntakaris et al., [n. d.]) is the most widely adopted LOB dataset within the field of deep learning applications to limit order books (Zhang et al., 2019, 2020; Tsantekidis et al., 2017a, b), particularly for forecasting endeavors. It comprises LOB data from five Finnish companies listed on the NASDAQ Nordic stock exchange: Kesko Oyj, Outokumpu Oyj, Sampo, Rautaruukki, and Wärtsilä Oyj. The data span ten trading days, from June 1st to June 14th, 2010, encompassing approximately 4 million limit order snapshots across ten levels of the LOB. The authors sampled LOB observations at intervals of ten events, resulting in a total of 394,337 samples. The label associated with each data point, indicative of mid-price movement, is determined by the percentage change between the prevailing mid-price:
(10) $p(t)=\frac{P^{ask}(t)+P^{bid}(t)}{2}$
and the average of the subsequent $h$ (chosen horizon) mid-prices:
(11) $m_{+}(t,k)=\frac{1}{k}\sum_{i=1}^{k}p(t+i)$
The percentage change is thus defined as:
(12) $l(t)=\frac{m_{+}(t,k)-p(t)}{p(t)}$
@@
where $k$ represents the window length, which in this instance also corresponds to the prediction horizon $h$. Labels are assigned as explained in 4. The dataset furnishes time series and corresponding class labels for five distinct horizons: $h\in H=\{10,20,30,50,100\}$. The dataset’s authors employed a uniform threshold $\theta=2\times 10^{-3}$ across all horizons. The value is chosen to balance the classes for $h=50$.
6.3. Experimental settings
@@
For each dataset, we trained and tested the performance of each model on different horizons, namely $10$, 2020, $50$, $100$. All the experiments were carried out using an RTX 3090. Since the FI-2010 dataset also contains 104 handcrafted features derived from the LOB, we used them in both our models. This choice improved the performance of the F1-Score (%) by approximately $1$. For Tesla and Intel, given the availability of message files containing the order information, we augmented the LOB snapshots by concatenating them with the corresponding orders. This integration was undertaken to incorporate additional information not present in the LOB. Consequently, this approach resulted in an approximate improvement of $1.5$ in the F1-score (%) . We report the details on the hyperparameters search in the Appendix (A).
@@
Baselines As comparative baselines, we employed 3 machine learning models: Support Vector Machine (SVM), Random Forest and XGBoost, and 8 deep learning SoTA LOB-based models: MLP, LSTM (Tsantekidis et al., 2017a), CNN (Tsantekidis et al., 2017b), CTABL (Tran et al., 2018), DAIN (Passalis et al., 2019), CNNLSTM (Tsantekidis et al., 2020), DeepLOB (Zhang et al., 2019) and BiN-CTABL (Tran et al., 2021). Due to computational constraints, we selected the top two performing models from FI-2010, specifically DeepLOB, and BiNCTABL, and exclusively trained and tested these models with the TSLA-INTC dataset.
@@
Trend Classification Threshold We remark that $\theta$ is the parameter that determines if a percentage change $l_{t}$ is classified as an up, stable, or downtrend. For the TSLA-INTC dataset, to ensure balanced class distribution, we set $\theta$ equal to the mean percentage change. In Sec. 6 we explore an alternative approach to defining $\theta$ based on financial parameters rather than class balance optimization. For the FI-2010 dataset, we retained the original labels to maintain consistency with existing benchmark studies and previous works.
@@
Metric We selected the F1-score as our primary performance metric because it captures both precision and recall in a single value. Accuracy is not a valid metric for our experiments because the classes are not balanced for each horizon. The F1-Score is robust to the class imbalance problem, which detrimentally affects the accuracy. Finally, the F1-score is the most used metric in the SoTA papers tackling the SPTP task. For a comprehensive evaluation, we provide precision and recall curves in the Appendix (B).
7. Results
@@
Table 2. F1-score on the FI-2010 dataset on four horizons. Bold values show the best scores.
@@
LSTM (Tsantekidis et al., 2017b) 66.5 58.8 66.9 59.4
CNN (Tsantekidis et al., 2017a) 49.3 46.1 65.8 67.2
CTABL (Tran et al., 2018) 69.5 62.4 71.6 73.9
DAIN-MLP (Passalis et al., 2019) 53.9 46.7 61.2 62.8
CNNLSTM (Tsantekidis et al., 2020) 63.5 49.1 69.2 71.0
DeepLOB (Zhang et al., 2019) 71.1 62.4 75.4 77.6
BiNCTABL (Tran et al., 2021) 81.1 71.5 87.7 92.1
@@
7.1. FI-2010 results
Table 2 presents the performance comparison across four prediction horizons [^5] Note that the horizon values represent the number of events before the sampling process of the dataset, while in the benchmarks (Prata et al., 2024; Ntakaris et al., [n. d.]) the values represent the horizons after the sampling process. In other words, the horizons considered are the same and are the ones defined originally in FI-2010. for the FI-2010 benchmark dataset. In the Appendix (B) we report also the precision and recall curves for horizon 100. MLPLOB and TLOB exhibit very high precision, also at high recall values, consistently achieve higher precision at all recall levels compared to the other models. The results for the baselines are extracted from the benchmark of Prata et al. (Prata et al., 2024)666if we had taken the results reported in the individual papers, MLPLOB and TLOB would have still outperformed all the other models. since the settings are equal for the FI-2010 dataset. MLPLOB and TLOB outperform all the other models analyzed in (Prata et al., 2024), surpassing state-of-the-art performance. Interestingly, MLPLOB demonstrates the best performance in the first three horizons. Notably, the performance differential between MLPLOB and TLOB is minimal, which, as we will demonstrate in Section 7.2, can be attributed to the lower complexity of the FI-2010 dataset, which explains the uselessness of a more complex architecture such as TLOB for this particular dataset.
7.2. Tesla and Intel results
Table 3. F1-score for Tesla on four horizons. Bold values show the best scores.

### v3

@@ 5.1 MLPLOB
$\mathbf{U}_{i,*}=\sigma\Bigl{(}\text{LayerNorm}\bigl{(}\sigma(\mathbf{X}_{i,*}\,\mathbf{W}_{1})\,\mathbf{W}_{2}\;+\;\mathbf{X}_{i,*}\bigr{)}\Bigr{)}\quad i=1,\dots,T,$ (8)
where $\sigma$ is the GeLU activation function [19], and LayerNorm denotes layer normalization.
@@
$\mathbf{Z}_{*,j}=\sigma\Bigl{(}\text{LayerNorm}\bigl{(}\sigma(\mathbf{U}_{*,j}\,\mathbf{W}_{3})\,\mathbf{W}_{4}\;+\;\mathbf{U}_{*,j}\bigr{)}\Bigr{)}\quad j=1,\dots,N.$ (9)
@@
Final Prediction. After several blocks of feature and temporal mixing, MLPLOB performs dimensionality reduction to collapse all features into a single vector, which then passes through several fully connected layers that gradually diminish the vector dimension and a final standard classification head. The network outputs the directional trend (up, down, or stable) for the final time step. Our primary objective in devising MLPLOB is to show that a carefully structured MLP-based model can match or exceed more complex architectures in the PTP task. The same method is also applied to TLOB.
5.2 TLOB
The Transformer architecture [41] has led to major breakthroughs in deep learning, notably in natural language processing [7, 22] and time-series modeling [42]. A key advantage is the ability to capture long-range dependencies without suffering as much from vanishing gradients or forgetting, and performance typically scales favorably with increased data [21]. Because massive volumes of financial LOB data are available and long-range dependencies are central in predicting the price trend, Transformers are well-positioned for LOB modeling.
@@ 5.2 TLOB
Temporal vs. Feature Attention. While standard Transformers [41] process tokens along a single dimension, LOB data naturally requires both temporal and spatial dependencies to be learned [34, 33]. For instance, time-step $t$ can reveal how deeper or shallower levels relate to one another, as well as how trends evolve over past snapshots. Hence, dual-attention explicitly addresses these two axes of variation. To investigate the importance of each type of attention layers we performed an ablation study (Section 7.6).
Bilinear Normalization Layer. To address non-stationarity and magnitude disparity (prices and sizes) in financial time series, we employ a Bilinear Normalization layer [37] as the initial layer. Unlike conventional $z$-score normalization, which can fail under distribution shifts at inference time, bilinear normalization adapts to batch-specific statistics, maintaining robust performance even when market conditions change. The same layer is also used in MLPLOB.
Positional Encoding. Because self-attention is permutation-invariant, we incorporate sinusoidal positional embeddings [41] to preserve the chronological structure within each LOB window. This embedding ensures that TLOB respects the temporal ordering of snapshots, which is crucial for modeling price evolution.
@@
6 Experiments
We conduct a comprehensive evaluation of MLPLOB and TLOB model training and testing on both the Benchmark FI-2010, the TSLA-INTC, and the BTC datasets. TLOB and MLPLOB surpass SoTA performances on every dataset and every horizon. TLOB performs the best on larger horizons, while MLPLOB performs the best on the shorter ones. Our experiments extend beyond merely demonstrating the state-of-the-art performance of TLOB, aiming to address several critical research questions: (1) Are stock prices harder to forecast than in the past? (2) What if we choose $\theta$ equal to the average spread? (3) Are temporal and spatial attention necessary? Through these investigations, we seek not only to validate our models’ predictive capabilities but also to contribute to the broader understanding of deep learning applications in financial forecasting.
Table 1: Intel and Tesla main characteristics for January 2015. Average liquidity is computed as the average quantity available in the first 10 LOB levels.
@@ 6 Experiments
TSLA $-0.42\pm 2.84$ $23,927,602\pm 4,554,884$ $0.16$ $3,320$
INTC $-0.44\pm 1.66$ $304,325,400\pm 69,340,430$ $0.01$ $124,960$
6.1 Benchmark dataset FI-2010
@@
The FI-2010 dataset [29] is the most widely adopted LOB dataset within the field of deep learning applications to limit order books [45, 46, 38, 39], particularly for forecasting endeavors. It comprises LOB data from five Finnish companies listed on the NASDAQ Nordic stock exchange: Kesko Oyj, Outokumpu Oyj, Sampo, Rautaruukki, and Wärtsilä Oyj. The data span ten trading days, from June 1st to June 14th, 2010, encompassing approximately 4 million limit order snapshots across ten levels of the LOB. The authors sampled LOB observations at intervals of ten events, resulting in a total of 394,337 samples. The dataset is pre-processed, with the labels already computed. The label associated with each data point, indicative of mid-price movement, is determined by the percentage change between the current mid-price and the average of the subsequent $h$ (chosen horizon) mid-prices. The percentage change is thus defined as:
$l(t)=\frac{m_{+}(t,k)-p(t)}{p(t)}$ (10)
@@
where $p(t)$ is the mid-price and $k$ represents the window length, which in this instance also corresponds to the prediction horizon $h$. Labels are assigned as explained in 4. The dataset furnishes time series and corresponding class labels for five distinct horizons: $h\in H=\{10,20,30,50,100\}$. The dataset’s authors employed a uniform threshold $\theta=2\times 10^{-3}$ across all horizons. The value is chosen to balance the classes for $h=50$.
6.2 BTC Dataset
The BTC dataset is extracted from Kaggle and encompasses 12 consecutive days, starting from January 9th, 2023, until January 20th, 2023. It comprises data from Binance Bitcoin perpetual data (BTCUSDT.P). The data points are already sampled at a frequency of 250 milliseconds, consequently, the four selected horizons (10, 20, 50, 100) correspond to a time horizon of 1s, 2.5s, 5s, and 10s, respectively. The dataset contains a total of 3.730.870 rows and it is partitioned such that the initial 9 days are allocated for training, the 10th day for validation, and the final day for testing.
6.3 TSLA-INTC Dataset
In the majority of state-of-the-art (SoTA) research within the domain of Deep Learning applied to LOB data, researchers typically employ one, two, or three stocks [10, 28, 26, 32, 20], predominantly from the technology sector. Adhering to this established practice, we construct a LOB dataset comprising two NASDAQ-listed stocks, namely, Tesla and Intel – spanning the period from January 2nd to January 30th, 2015. We posit that stylized facts and market microstructure characteristics exhibit independence from individual stock behaviors (as demonstrated in [6, 5, 14, 17] [^3] These seminal works in finance elucidate the universal statistical properties of LOBs, transcending specific stocks and markets.). The dataset encompasses 20 order book files for each stock, corresponding to each trading day, resulting in a total of approximately 24 million samples. The dataset is partitioned such that the initial 17 days are allocated for training, the 18th day for validation, and the final two days for testing. As shown in Table 1, the main characteristics of Tesla and Intel for January 2015 are very different, offering different market situations. Differently from TSLA, INTC is a small tick stock. Unfortunately, we cannot make the dataset public for copyright reasons.
Sampling. Limit Order Book data, especially for liquid stocks, is massive, every day, hundreds of thousands of orders are placed for each stock. Furthermore, financial data are known to have a low signal-to-noise ratio [27]. Accordingly, it is unnecessary to consider every LOB update, so defining a valid sampling technique is essential. While time-based [^4] BTC dataset is sampled every 100ms. and event-based sampling methods [^5] FI-2010 is sampled every 10 events. are used, they fail to capture the varying impact of transactions. In fact, single transactions can have very different impacts on the market. Volume-based sampling offers a solution by sampling the LOB after a predetermined volume of shares has been traded, thus reflecting the magnitude of market activity. Therefore, we adopted a sampling strategy based on trading volume, where snapshots of the Limit Order Book (LOB) are taken every 500 stocks traded. This method achieves a compromise between maintaining adequate temporal consistency within windows and ensuring significant variation between samples.
6.4 Experimental settings
@@
For each dataset, we trained and tested the performance of each model on different horizons, namely $10$, $20$, $50$, and 100100100100666depending on the sampling method, the horizon is in a different unit of measure. All the experiments were carried out using an RTX 3090. Since the FI-2010 dataset also contains 104 handcrafted features derived from the LOB, we used them in both our models. This choice improved the performance of the F1-Score by approximately $1$. For Tesla and Intel, given the availability of message files containing the order information, we augmented the LOB snapshots by concatenating them with the corresponding orders. This integration was undertaken to incorporate additional information not present in the LOB. Consequently, this approach resulted in an approximate improvement of $1.5$ in the F1-score. We report the details on the hyperparameter search in the Supplementary Material.
@@
Baselines As comparative baselines, we employed 3 machine learning models: Support Vector Machine (SVM), Random Forest and XGBoost, and 10 deep learning SoTA LOB-based models: MLP, LSTM [38], CNN [39], CTABL [36], DAIN [30], CNNLSTM [40], DeepLOB [45], BiN-CTABL [37], AXIALLOB [23] and DLA [18]. Due to computational constraints, we selected the top two performing models from [31], specifically DeepLOB and BiNCTABL, and exclusively trained and tested these models with the TSLA-INTC and BTC datasets. In Table 2, we report the inference time (ms) and the number of parameters for each SoTA model. Although TLOB and MLPLOB have a higher number of parameters compared to SoTA LOB-based models, they still have significantly fewer parameters than state-of-the-art
Table 2: Number of parameters and inference time for each model used in the experiments.
Model Nr. parameters Inference Time (ms)
MLP $10^{6}$ 0.08
LSTM [39] $1.6\cdot 10^{4}$ 0.21
CNN [38] $3.5\cdot 10^{4}$ 0.36
CTABL [36] $1.1\cdot 10^{4}$ 0.48
DAIN-MLP [30] $5.3\cdot 10^{4}$ 0.50
CNNLSTM [40] $2.8\cdot 10^{5}$ 0.49
AXIALLOB [23] $2\cdot 10^{4}$ 1.91
DLA [18] $1.2\cdot 10^{5}$ 0.23
DeepLOB [45] $1.4\cdot 10^{5}$ 1.31
BiNCTABL [37] $1.1\cdot 10^{4}$ 0.71
MLPLOB $6.3\cdot 10^{7}$ 4.79
TLOB $1\cdot 10^{7}$ 2.24
deep learning models commonly used in standard machine learning tasks. Furthermore, they do result in slightly higher inference times, but their speed remains adequate for application in high-frequency trading scenarios.
@@
Trend Classification Threshold We remark that $\theta$ is the parameter that determines if a percentage change $l_{t}$ is classified as an up, stable, or downtrend. For the TSLA-INTC and the BTC datasets, to ensure balanced class distribution, we set $\theta$ equal to the mean percentage change. In Sec. 8 we explore an alternative approach to defining $\theta$ based on financial parameters rather than class balance optimization. For the FI-2010 dataset, we retained the original labels to maintain consistency with existing benchmark studies and previous works.
@@
Metric We selected the F1-score as our primary performance metric because it captures both precision and recall in a single value. Accuracy is not a valid metric for our experiments because the classes are not balanced for each horizon. The F1-score is robust to the class imbalance problem, which detrimentally affects the accuracy. Finally, the F1-score is the most used metric in the SoTA papers tackling the PTP task. For a comprehensive evaluation, we provide precision and recall curves in the Supplementary Material.
7 Results
@@
Table 3: F1-score on the FI-2010 dataset on four horizons. Bold values show the best scores.
@@
LSTM [39] 66.5 58.8 66.9 59.4
CNN [38] 49.3 46.1 65.8 67.2
CTABL [36] 69.5 62.4 71.6 73.9
DAIN-MLP [30] 53.9 46.7 61.2 62.8
CNNLSTM [40] 63.5 49.1 69.2 71.0
AXIALLOB [23] 73.2 63.4 78.3 79.2
DLA [18] 79.4 69.3 87.1 52.2
DeepLOB [45] 71.1 62.4 75.4 77.6
BiNCTABL [37] 81.1 71.5 87.7 92.1
@@
7.1 FI-2010 results
Table 3 presents the performance comparison across four prediction horizons [^7] Note that the horizon values represent the number of events before the sampling process of the dataset, while in the benchmarks [31, 29] the values represent the horizons after the sampling process. In other words, the horizons considered are the same and are the ones defined originally in FI-2010. for the FI-2010 benchmark dataset. In the Supplementary Material, we also report the precision and recall curves for horizon 100. MLPLOB and TLOB exhibit very high precision, also at high recall values, and consistently achieve higher precision at all recall levels compared to the other models. The results for the baselines are extracted from the benchmark of Prata et al. [31]888if we had taken the results reported in the individual papers, MLPLOB and TLOB would have still outperformed all the other models. since the settings are equal for the FI-2010 dataset. MLPLOB and TLOB outperform all the SoTA LOB-based models, surpassing state-of-the-art performance. Notably, the performance differential between MLPLOB and TLOB is minimal, which, as we will demonstrate in Section 7.2, can be attributed to the lower complexity of the FI-2010 dataset, which explains the uselessness of a more complex architecture such as TLOB for this particular dataset.
7.2 Tesla and Intel results
Table 4: F1-score for Tesla on four horizons. Bold values show the best scores.